OpenStudy (anonymous):

whats the limit t-->0^+ of (9t^t)?

OpenStudy (anonymous):

is this \[\lim_{t\rightarrow 0^+} 9t^t\]?

OpenStudy (anonymous):

can take the 9 right outside because it is a constant and compute \[9\lim_{t\rightarrow 0^+} t^t\]

OpenStudy (anonymous):

to compute this limit, you have a choice. either "take the log" or rewrite \[t^t=e^{t\ln(t)}\] but all the work is the same you were probably told to "take the log"

OpenStudy (anonymous):

yeah I know it's l'hopitals rule but I can't get it o work

OpenStudy (anonymous):

yeah take the log

OpenStudy (anonymous):

ok so first step is take the log and get \[t\ln(t)\] right?

OpenStudy (anonymous):

then it is of the form \[0\times -\infty\] so you use the standard trick of rewriting it as \[\frac{\ln(t)}{\frac{1}{t}}\] and now it looks like \[\frac{\infty}{\infty}\] so you can use l'hopital

OpenStudy (anonymous):

the derivative of the log is \[\frac{1}{t}\] and the derivative of \[\frac{1}{t}\] is \[-\frac{1}{t^2}\]

OpenStudy (anonymous):

but if you put in 0 it's still infinity?

OpenStudy (anonymous):

so after l'hopital you get \[\frac{\frac{1}{t}}{-\frac{1}{t^2}}=-t\]

OpenStudy (anonymous):

now the limit is clearly 0, so your original limit is \[9e^0=9\]

OpenStudy (anonymous):

Oh I see that make sense thanks a lot for your help

OpenStudy (anonymous):

clear that we have to "exponentiate" right because the first step was to "take the log" so we computed the limit of the log, not the limit we were after

OpenStudy (anonymous):

yw, and this is a pretty common limit, so you might just remember that \[\lim_{x\rightarrow 0^+}x^x=1\]